OpenStudy (anonymous):

Find the scalar equation of the plane containing the points A(-3, 1, 1) and B(-4, 0, 3) and the vector u = [1, 2, 3] Are my steps right? Make one a point a vector. Cross the new vector with u. Then put it in scalar form with the remaining point?

OpenStudy (amistre64):

you cant simply make a point a vector. a vector is defined by 2 points

OpenStudy (amistre64):

the vector from B to A is: A(-3, 1, 1) -B(-4, 0, 3) ------------ [1, 1,-2 ] now you can cross that

OpenStudy (amistre64):

once you have your crossed (normal vector) you can then attach it to a given point

OpenStudy (anonymous):

Yep that's what I meant. 7x+5y+11z-5 was my final answer. Is that correct?

OpenStudy (amistre64):

x 1 1 y 2 1 z 3-2 x = -4-3 y = -(-2-3) z = 1-2 [7, -5, 1] unless i mis crossed; i would have to disagree with your results

OpenStudy (amistre64):

7x - 5y + z + d = 0 such that d is the correct constant from the expansion stuff